OpenStudy (anonymous):

A 20−pound bag of sugar costs (y + 3) dollars while a 10−pound package costs (y − 3) dollars. If the price per pound of sugar is the same for both packages, the equation below can be used to solve for y. y+3/20 = y-3/10 What is the value, in dollars, of y?

OpenStudy (ash2326):

We have \[\frac{y+3}{20}=\frac{y-3}{10}\] multiply both sides by 20, can you do it @rebecca1233 ?

OpenStudy (anonymous):

which side 10*20?

OpenStudy (ash2326):

\[20\times \frac{y+3}{20}=\frac{y-3}{10}\times 20\]

OpenStudy (ash2326):

could you simplify this?

OpenStudy (anonymous):

oh alright, give me a sec (:

OpenStudy (anonymous):

i think so

OpenStudy (ash2326):

cool :)

OpenStudy (anonymous):

i got 9 am i correct :)

OpenStudy (ash2326):

yes, good work:D did you understand?

OpenStudy (anonymous):

yes :)